OpenStudy (anonymous):

if x-y=7 and xy=-23/2. find x ^2+y^2. please show me the steps.

OpenStudy (anonymous):

\[(x-y)^2 = 7^2\]\[x^2-2xy+y^2=49\] plug in the value of xy

OpenStudy (anonymous):

dam son you smart

OpenStudy (anonymous):

I didn't get it .Please show me more steps

OpenStudy (anonymous):

Given that \[x - y = 7\text{ and }xy = -\frac{23}{2}\]squaring both sides yields\[(x-y)^2 = 7^2\]\[x^2 - 2xy + y^2 = 49\]now substitute xy with its value\[x^2 - 2(-\frac{23}{2})+y^2 = 49\]\[x^2 + 23 + y^2 = 49\]\[x^2 + y^2= 26\]

OpenStudy (anonymous):

Thank you very much